### v1

# Respiratory Subtraction for Pulmonary Microwave Ablation Evaluation

Abstract
Currently, lung cancer is a leading cause of global cancer mortality, often necessitating minimally invasive interventions. Microwave ablation (MWA) is extensively utilized for both primary and secondary lung tumors. Although numerous clinical guidelines and standards for MWA have been established, the clinical evaluation of ablation surgery remains challenging and requires long-term patient follow-up for confirmation. In this paper, we propose a method termed respiratory subtraction to evaluate lung tumor ablation therapy performance based on pre- and post-operative image guidance. Initially, preoperative images undergo coarse rigid registration to their corresponding postoperative positions, followed by further non-rigid registration. Subsequently, subtraction images are generated by subtracting the registered preoperative images from the postoperative ones. Furthermore, to enhance the clinical assessment of MWA treatment performance, we devise a quantitative analysis metric to evaluate ablation efficacy by comparing differences between tumor areas and treatment areas. To the best of our knowledge, this is the pioneering work in the field to facilitate the assessment of MWA surgery performance on pulmonary tumors. Extensive experiments involving 35 clinical cases further validate the efficacy of the respiratory subtraction method. The experimental results confirm the effectiveness of the respiratory subtraction method and the proposed quantitative evaluation metric in assessing lung tumor treatment.
I Introduction
Lung cancer stands as a formidable global health challenge, with consistent statistics from the World Health Organization (WHO) indicating its prominence as the leading cause of cancer-related mortality worldwide. In 2020 alone, an estimated 2.2 million new cases of lung cancer and 1.8 million associated deaths were reported[1]. The defining characteristic of lung cancer lies in the development of malignant tumors that arise from diverse cell types and tissues within the pulmonary structure[2]. According to the guidelines set forth by the National Comprehensive Cancer Network (NCCN), surgical resection is considered the primary treatment modality, involving the excision of both the tumor and surrounding lung tissue[3]. Nevertheless, this procedure often entails the removal of considerable amounts of healthy lung tissue, rendering it less cost-effective, particularly for addressing small tumors in their initial stages [4, 5]. Furthermore, a considerable portion of patients encounter barriers to undergoing surgical resection, owing to factors such as age, physical conditions, or reluctance to undergo the procedure [6]. Recently, the utilization of image-guided thermal ablation techniques, including radio-frequency ablation (RFA) and microwave ablation (MWA), has increased, primarily due to reduced invasiveness[7, 8, 9]. Compared to RFA, MWA offers better control over the ablation range in the pulmonary environment due to the lungs’ low electrical conductivity. Extensive research on pulmonary MWA has provided substantial data supporting its efficacy and safety in treating lung tumors[10, 11, 12, 13]. As a result, MWA has become a preferred therapeutic option for patients who are not candidates for surgical resection or are medically inoperable. This procedure involves the percutaneous insertion of a needle-like ablation probe, guided by CT scans, into the tumor site. Once positioned correctly, the probe’s tip thermally destroys the cancer cells in its vicinity.
Although MWA offers a precisely targeted coverage area reducing damage to normal tissue, it potentially increases the risk of residual tumors. This is primarily because doctors cannot rigorously assess whether the MWA ablation covers the entire tumor area and whether the tumor has been completely ablated. After MWA treatment, an additional CT scan is typically performed to compare with preoperative images, allowing for visual assessment of the treatment’s efficacy. However, it often fails to provide doctors with sufficient direct judgment criteria. Due to lung deformation caused by necrotic cancer cells during ablation, structural differences between preoperative and postoperative images make it difficult to directly link the preoperative tumor mask with the postoperative image[8]. Additionally, there are highlighted areas in the postoperative image due to hemorrhage, tissue necrosis, etc., which interfere with residual tumor locating. Also, the correlation between appropriate ablation areas and tumor sizes in lung ablation surgeries has been extensively studied to guide surgical procedures[14]. However, there remains a lack of comprehensive research on methods to assess the efficacy of ablation surgeries without long-term follow-up.
Figure 1: Overall framework for respiratory subtraction quantitative assessment
To better evaluate MWA results and assist in clinical decision-making, this paper introduces a novel assessment method called "respiratory subtraction." First, the tumor area in preoperative images and the treatment areas in postoperative images are segmented. Then, preoperative images are registered to postoperative images to unify respiratory phases using a coarse rigid registration followed by a non-rigid registration. Finally, respiratory subtraction images are obtained by subtracting the registered preoperative images from the postoperative images. Doctors can assess the completeness of the ablation by observing the coverage area of the treatment region in the subtraction images, thereby enhancing the success rate of the therapy and reducing the risk of tumor recurrence due to incomplete ablation. Additionally, we propose a novel quantitative evaluation metric, the Ablation Effectiveness Scale (AES), to assess the treatment coverage of tumors. Experimental results demonstrate the exceptional efficacy of the proposed respiratory subtraction method for evaluating the quality of MWA surgery. Our method offers a more direct and effective method for evaluating clinical surgical efficacy compared to the current approach by which doctors rely on visual observation of both the preoperative and postoperative images and long-term follow-up records.
II Method
The overall framework of our method of the pulmonary MWA respiratory subtraction is shown in Fig. 1. In general, the lung parenchymas are segmented and registered from both preoperative and postoperative images. The respiratory subtraction images are then acquired by subtraction. The subtraction images are subsequently used in qualitative and quantitative assessments of surgical effectiveness.
II-A Lung Parenchyma Segmentation
For processing pulmonary MWA data, registration to reduce respiratory motion is accomplished by separately extracting the lung parenchyma. This approach is adopted because the clinical focus is solely on the lung parenchyma, and it improves the segmentation performance of tumors and treatment areas, leading to more accurate assessments. Specifically, Unet-R231[15], an advanced structure based on Unet known for its superior lung segmentation capabilities, is the first step of our method. This model generates lung parenchyma masks for both preoperative and postoperative images. Trained on a clinical routine dataset (n=231) containing multiple diseases, the model provides masked images, $P_{pre}$ and $P_{post}$. These masked images serve as the moving and fixed images, respectively, for subsequent rigid registration of the lung parenchyma.
II-B Lung Parenchyma Rigid Registration
The initial coarse registration of lung parenchyma is then performed to ensure the general placement and alignment of the preoperative and postoperative images. The three-dimensional rigid registration is conducted on the extracted preoperative and postoperative lung parenchyma images $P_{pre}$ and $P_{post}$ by applying rotations and translations on the moving image $P_{pre}$. Aiming to minimize the differences between the $P_{pre}$ and $P_{post}$, the alignment process is quantified using a loss function normalized cross-correlation (NCC). NCC measures the similarity between the two images and is defined as:
$\text{NCC}(P_{\text{pre}},P_{\text{post}})=\\ \frac{\sum_{i,j,k}\left(P_{\text{pre}}(i,j,k)-\bar{P}_{\text{pre}}\right)\left(P_{\text{post}}(i,j,k)-\bar{P}_{\text{post}}\right)}{\sqrt{\sum_{i,j,k}\left(P_{\text{pre}}(i,j,k)-\bar{P}_{\text{pre}}\right)^{2}\sum_{i,j,k}\left(P_{\text{post}}(i,j,k)-\bar{P}_{\text{post}}\right)^{2}}}$
where $\bar{P}_{\text{pre}}$ and $\bar{P}_{\text{post}}$ are the mean intensities of the preoperative and postoperative images, respectively. The goal of the rigid registration is to find the transformation $T_{1}$ that maximizes the NCC value, indicating the best alignment of the images.
The resulting rigid transformation $T_{1}$ can be mathematically described as:
$T_{1}(\mathbf{x})=R\mathbf{x}+\mathbf{t}$
where $R$ is a $3\times 3$ rotation matrix, $\mathbf{t}$ is a translation vector, and $\mathbf{x}$ represents the coordinates of the points in the image. The transformation $T_{1}$ is optimized iteratively to align $P_{\text{pre}}$ with $P_{\text{post}}$. This optimization is performed using algorithms available in the SimpleITK library, which adjust the parameters of $R$ and $\mathbf{t}$ to improve the similarity measure given by the NCC. After this step, the lung parenchyma from the preoperative images is adjusted to the position of the lung parenchyma from the postoperative images, preparing for the non-rigid registration of lung tissues.
II-C Tumor Segmentation and Treatment Area Segmentation
Before non-rigid registration of lung tissues, it is necessary in preprocessing to segment and remove tumors from the preoperative images and treatment areas from the postoperative images. Retaining these regions could cause them to be interpreted as prominent features, adversely affecting the non-rigid registration process. To minimize such interference, tumor region masks $M_{pre}$ are extracted from the preoperative images $P_{pre}$, and treatment area masks $M_{post}$ are extracted from the postoperative images $P_{post}$. These masks are automatically generated using a Unet-based architecture trained by the teacher-student approach[16] and then the corresponding part is excluded from the images.
II-D DeedsBCV Non-Rigid Registration
Considering the characteristics of lung tissues and the possible significant deformations in lung areas between preoperative and postoperative stages, a non-rigid registration method, deedsBCV[17], is introduced to further fine-tune the registration results. Deeds in deedsBCV stands for DEnsE Displacement Sampling. It applies a hierarchical, multi-resolution approach to calculate dense displacement fields, gradually optimizing at different resolution levels to handle both large-scale and small-scale deformations in the images. This method constructs a graph model using control points in a uniform B-spline grid as nodes, with an optimized energy function composed of a data term and regularization term:
$E(f)=\underbrace{\sum_{p\in P}D(f_{p})}_{\text{data term}}+\alpha\underbrace{\sum_{(p,q)\in N}R(f_{p},f_{q})}_{\text{regularization term}}.$ (1)
where $p$ denotes the control points corresponding to the spatial location $x_{p}$, and for each node, $f_{p}$ is the discrete displacements. The node $q$ denotes the direct neighbor node connected to $p$. $\alpha$ is a hyperparameter that controls the balance between the data term and the regularization term. The data term describes the similarity between the control point in one image and the corresponding control point in another image with discrete displacement $f_{p}$. In our implementation, the self-similarity descriptor is adopted to calculate the similarity of the corresponding nodes. The pairwise regularization term ensures global smooth transformation by penalizing the squared displacement differences of adjacent control points:
$R(f_{p},f_{q})=\sum_{(p,q)\in{N}}\frac{\left\|\mathbf{u}_{p}-\mathbf{u}_{q}\right\|^{2}}{\left\|\mathbf{x}_{p}-\mathbf{x}_{q}\right\|}$ (2)
A great advantage of deedsBCV is its discrete optimization framework. Unlike traditional gradient-based optimization methods, deedsBCV operates within a discrete label space. Each voxel in the image is assigned a discrete displacement label, representing potential deformations at that point. This discrete approach helps to avoid common local optima found in continuous optimization, thereby achieving more accurate registration results.
After non-rigid registration, we get the non-rigid deformation field $T_{2}$. The final registered preoperative image $P_{registered}$ can be obtained using the equation
$P_{registered}=P_{pre}\circ T_{1}\circ T_{2}$ (3)
where $\circ$ represents the registration operation.
II-E Subtraction
After aligning the preoperative and postoperative images, the respiratory subtraction images $P_{sub}$ can be easily obtained by subtracting the registered preoperative images from the postoperative images:
$P_{sub}=P_{post}-P_{registered}=P_{post}-P_{pre}\circ T_{1}\circ T_{2}$ (4)
The resulting subtraction image $P_{sub}$ offers enhanced visual confirmation of lung MWA treatment performance compared to traditional observation methods, and visual assessment can be made by comparing the tumor area and treatment boundary in $P_{sub}$. Additionally, quantitative evaluation can be achieved and will be introduced in our next subsection.
II-F Treatment Evaluation
To further assist doctors in assessing surgery, in addition to visualization aid provided by the subtraction image $P_{sub}$, we propose a new quantitative metric based on $P_{sub}$ to evaluate the lung MWA treatment effectiveness, named Ablation Effectiveness Scale (AES) index.
An effective treatment should include the tumor area within the treatment area to ensure complete tumor ablation, but it should not extend excessively beyond which may cause damage to surrounding healthy tissues. The treatment boundary should aim to minimize the treatment area while still achieving complete tumor ablation. Clinically, it is generally recommended that the ablation range extends 5-10mm beyond the tumor border to ensure the edges of the tumor are treated and to reduce the risk of residual lesions.
Figure 2: Respiratory subtraction ablation therapy evaluation (a) Under-ablation; (b) Average-ablation; (c) Over-ablation; $T$: Tumor; $B$: Actual ablation; $A$: Ideal ablation
In clinical scenarios, the situation of ablation surgery can be broadly categorized into three types: under-ablation, average-ablation (or perfect ablation), and over-ablation (or excessive ablation), as shown in Fig.2. Under-ablation refers to the situation where the tumor area is not completely covered by the ablation area, indicating a surgical failure and potentially necessitating an additional ablation surgery for the patient. Average-ablation indicates that the tumor area is completely covered by the ablation area, and the ablation area is within the optimal range for the procedure. Over-ablation means that while the tumor area is completely covered, the ablation area exceeds the clinically acceptable range, potentially causing unnecessary damage to normal tissues.
To better reflect the therapeutic outcomes of ablation surgery, our metric, the AES index, is designed to assess potential scenarios in lung ablation surgery from three aspects: actual treatment tumor coverage, ideal ablation area coverage, and the degree of over-ablation. To facilitate understanding, the symbols and their meanings in the subsequent formulas are defined as follows. The tumor volume and ablation volume are denoted as $T$ and $B$, respectively. The ideal ablation volume of the tumor is denoted as $A$, which is obtained by expanding the boundary of volume T by 5mm outward. The symbols $\cap$ and $\setminus$ represent the intersection and difference of two sets, respectively. Based on these definitions, the AES index can be calculated using Eq 8 to evaluate the effectiveness of lung MWA surgery:
$CR_{1}=\frac{T\cap B}{T}$ (5)
$CR_{2}=\frac{B\cap A}{A}$ (6)
$ER=\frac{B\setminus A}{B}$ (7)
${AES=\left\{\begin{matrix}-1+e^{-\alpha\cdot(1-CR_{1})-\beta\cdot ER}&if(CR_{1}<\lambda)\par\\ 1-e^{-\gamma\cdot(1-CR_{2})-\theta\cdot ER}&if(CR_{1}\geq\lambda)\end{matrix}\right.}$ (8)
where $CR_{1}$ denotes the current tumor coverage ratio of the ablation area, $CR_{2}$ is the ratio of the current valid ablation area to the ideal ablation area and $ER$ is used to represent the degree of over-ablation. The AES index is specifically defined in Eq 8. In the case of low tumor coverage ($CR_{1}$ < $\lambda$), the primary focus of the AES index lies in ensuring complete tumor coverage and evaluating the proportion of invalid ablation area. Empirically, hyperparameters $\alpha$ and $\beta$ are set to 1.0 and 2.0, respectively, to balance tumor coverage and over-ablation rate in subsequent experiments. As tumor coverage increases in this scenario, the AES index value also increases gradually towards zero. While at a similar level of tumor coverage, a higher degree of over-ablation results in a lower AES index value. When tumor coverage exceeds threshold $\lambda$, the AES index primarily emphasizes assessing ablation area coverage and over-ablation proportion. Similarly, hyperparameters $\gamma$ and $\theta$ are empirically set to 0.5 and 2.0 respectively to strike a balance between these factors. In the second scenario, as ablation area coverage increases, there is a decrease in the corresponding AES index value. While at an equivalent level of ablation area coverage, less over-ablation leads to smaller values approaching zero. In our experiment, the parameter $\lambda$ is empirically set to 0.75.
III Experiment and results
III-A Dataset
In this study, clinical preoperative and postoperative CT images were collected at the time of immediate postoperative, 1 month post-surgery, 3 months post-surgery, and 6 months post-surgery from 35 patients at Beijing Chaoyang Hospital, and written informed consent was provided by all patients. In our experiment, all the involved masks were manually annotated by doctors. All the CT data underwent preprocessing, which included cropping to the dimension of 256*256*256 and configuring the spacing to 1.25mm*1.25mm*1.25mm. Subsequently, normalization was performed to scale the pixel values of the data within the range between zero and one. The proposed respiratory subtraction method and the aforementioned evaluation metrics were utilized in the conducted experiments. For all statistical analyses, SPSS software was used to perform Kappa analysis and Spearman correlation analysis.
Figure 3: HU distribution comparison between the preoperative tumor area and postoperative ablation treatment area
Figure 4: Registration results after tumor segmentation and corresponding HSV pseudo-color images on the real clinical data. The HSV pseudo-color images represent the difference between the registered moved images and the original postoperative images. Preoperative tumor masks after registration are remapped onto the HSV images and displayed in yellow, while the postoperative treatment area boundaries are depicted in blue. Red regions in these HSV images represent significant changes in HU values within the ablation area before and after the procedure. The image depicts a respiratory subtraction example of an unsuccessful ablation procedure of a patient, which has been confirmed by medical records that this patient underwent surgical intervention for the tumor again after three months. It can be evidently observed that the ablation area (blue) did not completely encompass the tumor (yellow).
III-B Statistical analysis of preoperative and postoperative HU around the tumor region
The Hounsfield Unit (HU) values around the tumor region were statistically analyzed both preoperative and postoperative, as depicted in Fig. 3. The analyzed region specifically refers to the area outside the tumor area but within the treatment area. The statistical analysis reveals a significant difference in HU values in the region before and after surgery, with the two distributions distinguishable using a threshold of around -600 HU. This statistical finding has significant clinical implications for evaluating lung ablation surgeries. It allows for determining the range of ablation by assessing the average HU value at the ablation site. The finding aligns with the clinical follow-up outcomes obtained from 60 patients.
III-C Registration results for respiratory subtraction
To demonstrate the efficacy of respiratory subtraction, we conducted experiments on the collected dataset and visualized the subtraction results as HSV pseudo-color images. as depicted in Fig.4. The figure illustrates that rigid registration can roughly align the images, while non-rigid registration precisely aligns the lung tissues in detail. By comparing the images in the "Moved (Non-Rigid)" column with those in the "Fixed" column, it is evident that the detailed lung tissue structures in the moved (preoperative) images are well-aligned with those in the fixed (postoperative) images. The findings indicate that the proposed respiratory subtraction method achieves highly accurate registration results in critical soft tissues of the lungs, demonstrating precise mapping of fundamental lung features to their corresponding positions.
The accuracy of the respiratory subtraction registration process was further quantitatively analyzed using normalized cross-correlation (NCC), structural similarity (SSIM), root mean square error (RMSE), and Dice coefficients for lung parenchyma masks derived from postoperative and registered preoperative images. Among them, rigid registration based on mutual information (MI) and FastSymmetricForcesDemonsRegistrationFilter (FSFDRF)[18] were selected for comparison with our registration algorithm. The rigid registration algorithm based on mutual information is remarkably effective in aligning multi-modal medical images, showcasing exceptional robustness and precision in clinical scenarios. In parallel, the FSFDRF algorithm excels in handling complex deformations and anatomical variances, making it suitable for scenarios such as lung ablation surgery.
Table I presents the results of these registration metrics on the experiment’s clinical data. The results demonstrate that our method achieved a significant improvement in Dice coefficients compared with both MI registration and FSFDRF registration, all with values exceeding 0.90. Among these methods, our algorithm achieved the highest Dice coefficient of 0.98, indicating superior registration performance. Furthermore, there was a notable decrease in RMSE, which also indicated that our registration process effectively aligns overall lung regions. Additionally, noteworthy enhancements were observed in NCC and SSIM, suggesting that our registration process effectively aligns local lung details to meet clinical application requirements and provide solid performance for clinical decision-making and evaluation in pulmonary MWA surgery.
TABLE I: Quantitative comparison of clinical data registration results of different registration methods
NCC $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ Dice $\uparrow$
No Registration 0.6301 0.8040 230.76 0.7015
MI Rigid Registration 0.8805 0.8567 131.04 0.9002
FSFDRF 0.9398 0.8985 89.09 0.9478
Our method 0.9733 0.9231 59.34 0.9791
III-D Validity analysis of AES index
Figure 5: The distribution of doctors’ subjective rating on images during and 1-6 months postoperative
Immediately\6 Months Under Average Over
Under ablation 7.14% 0.00% 0.00%
Average ablation 0.00% 0.00% 0.00%
Over ablation 28.57% 50.00% 14.29%
TABLE II: Confusion matrix between doctor-ratings immediately postoperative and 6 months postoperative
To validate the effectiveness of the AES index in evaluating ablation surgery efficacy, two highly skilled doctors specializing in lung ablation surgery were invited to independently rate subjective scores for all patients’ CT images at different time points: immediately postoperative, 1 month postoperative, 3 months postoperative, and 6 months postoperative. The ratings of CT images taken immediately after surgery and during follow-up (paired with all medical records) were used to compare the correlation with our AES index metric. Additional postoperative follow-up medical records, such as images showing proliferating tumors when initial ablation surgery fails, were also considered. The doctors assigned subjective scores categorized into three levels corresponding to common ablation surgery outcomes: 1 for over-ablation, 0 for average ablation, and -1 for under-ablation. The collected scores demonstrated a high Kappa value of 0.645 (p<0.001), indicating a high level of consistency in the doctors’ assessments.
Figure 6: Illustration of the two special cases. (a1) and (b1) represent the preoperative CT images of two cases, with the tumor areas indicated in red. (a2) and (b2) represent the corresponding postoperative CT images, where the green areas denote the treatment regions and the red areas represent the registered tumor areas.
Notably, our experiments identified several outlier data points among cases classified as over-ablation (AES index value of 1) by doctors on immediate postoperative images, our system predicted an AES index value significantly less than 0, approaching -1, indicating potential under-ablation and possible therapeutic failure. Two such cases are illustrated in Fig. 6. Both cases have small tumor areas and large ablation areas, which led doctors to believe the ablation surgery was successful. However, As observed in (a2) and (b2) of Fig. 6, the registered tumor areas (red) are not included in the large ablation areas (green), indicating under-ablation. This is also objectively confirmed by medical records, which revealed that one patient was identified as under-ablated during the one-month follow-up, while another patient was identified as under-ablated during the six-month follow-up and subsequently underwent another ablation procedure.
Fig. 5 illustrates the class distribution between doctors’ assessments. The longer-term postoperative ratings paired with medical records can be considered as the ground truth assessment since it was based on all available information and the subsequent results of the therapy. The over-proportion of the over-ablation labels in the left part of the figure shows that, on immediate postoperative images, doctors tend to overestimate the ablation ratings and many cases that should be labeled as under-ablation or average-ablation were miss-labeled as over-ablation. The possible reason is that by just looking at the immediate postoperative image, doctors may not be able to remember the exact location of the preoperative cancer area due to lung deformations, therefore they may tend to assess the treatment majorly based on the absolute ablation area. Table II shows the confusion matrices between the doctors’ subjective scores immediately after surgery and 6 months after. In Table IV, we can find that more than 78% of cases judged as over-ablation by doctors immediately after surgery were later on rated as under-ablation or average-ablation. This proves that doctors tend to overestimate the ablation ratings when merely judging from the images immediately after surgery.
This phenomenon suggests that knowing the exact location of the original cancer in postoperative images is vital in preventing misjudgments. Our method provides both image and score to assist the doctors so that they can better understand the relative position between the postoperative ablation area and the preoperative cancer area immediately after surgery. These cases also prove the importance of our registration system which can provide the exact location of cancer even after it disappeared after the ablation. Our respiratory subtraction image could give doctors a solid referencing point when assessing whether the ablation treatment is sufficient and therefore prevent misjudgment.
As shown in Table III, the Spearman correlation analysis results indicate a robust positive correlation between the AES index and the doctors’ ratings (6 months after surgery), which suggests that the AES index can provide valuable support in assessing clinical ablation procedures. Compared with the Spearman correlation coefficient value of 0.443 from the immediate postoperative result, the value of 0.788 indicates that our method shows a stronger correlation with the accurate long-term judgment results. Meanwhile, Table IV shows a confusion matrix between the doctors’ subjective scores 6 months after and our AES index (thresholded by $\lambda$). From table IV we can see that our method can predict the under-ablation cases with full precision, with all the under-ablation cases in the experiment dataset accurately found and rated. As mentioned, under-ablation scenarios can potentially cause dangerous mistreatment or readmission to excessive therapies afterwards. This result shows that our method could preemptively prevent such possibilities caused by doctors’ overlook of under-ablation scenarios and protect patients from missing the best treatment opportunity or being subjected to unnecessary second ablation therapy afterwards.
TABLE III: Correlation between AES index (immediately postoperative) and doctor ratings (immediately postoperative and 6 months postoperative respectively)
Metric Immediately 6 Months After
Spearman Correlation 0.443 0.788
P-value 0.0077 0.0006
TABLE IV: Confusion matrix between 6-month postoperative doctor-rating and AES index
[…]
Under ablation 21.43% 0.00% 0.00%
Average ablation 0.00% 21.43% 0.00%
Over ablation 0.00% 28.57% 28.57%
IV Conclusion and Discussion
In this study, we proposed the respiratory subtraction method and the AES metric for lung tumor ablation therapy performance evaluation, and collected a dataset of 35 cases for preliminary validation of the algorithm and metric. Based on the ratings provided by doctors in Fig. 5, it is evident that the majority of the immediate postoperative data tended to be labeled as over-ablation by doctors, while only about 1/5 of the under-ablation cases are correctly discovered (see Table II). Relying solely on visual inspection poses challenges for doctors in ensuring complete ablation. Consequently, to maximize therapeutic success and improve patient survival rates, doctors tend to employ over-ablation methods as a precautionary measure. This situation highlights the current lack of effective evaluation methods for clinical lung ablation therapy and the need for accurate aiding tools in ablation effectiveness judgement for doctors.
Moreover, delineating the preoperative tumor overlay and the ablation area on immediate postoperative images is challenging. Doctors typically need to complete real-time assessments immediately after treatment while the lung deformation is significant between preoperative and postoperative images. Hence, there is an urgent need to develop an automated system to offer assistance. The imaging-guided respiratory subtraction system we developed for evaluating the performance of lung MWA treatment can automatically provide doctors with both accurate preoperative tumor overlays on the deformed postoperative treatment region and numerical metrics to evaluate treatment efficacy. Judging from the results shown in Tab. III and Tab. IV, our methods can successfully discover under-ablation scenarios and achieve accuracy in long-term clinical outcomes through a streamlined automatic process. When the majority of the cases viewed by doctors immediately after surgery tend to be categorized as over-ablation (see Tab. II), our method can serve as a supplemented reminder for clinical decisions in order to avoid unnecessary harm to patients.
In our system, preoperative images are aligned with postoperative images using both coarse rigid and the deedsBCV non-rigid registration method. Subsequently, the registered preoperative images are subtracted from the postoperative images to obtain the respiratory subtraction images. To enhance the clinical assessment of MWA treatment, we introduce a quantitative analysis scheme to evaluate discrepancies between preoperative tumor regions and postoperative treatment coverage. Experimental results using clinical data confirm the effectiveness of both the respiratory subtraction images and our proposed criterion AES index for assessing the ablation area effectiveness during lung tumor ablation surgery.
Moving forward, we plan to establish a multi-center dataset to collect more diverse data and validate metric effectiveness across various clinical scenarios. Additionally, this study calculates and evaluates respiratory subtraction and ablation area effectiveness metrics based on preoperative and postoperative CT images. To enhance the clinical utility of respiratory subtraction methodology further, we will develop a large-scale preoperative CT/ intraoperative CBCT dataset and apply this technology during intraoperative settings enabling real-time insights into ablative processes and facilitating effective judgment regarding surgery progress while allowing adjustments in instances involving under-ablations or minimizing severe over-ablation without compromising surrounding healthy tissues.